OpenStudy (anonymous):

find reciprocal of -9b over 6b+2

OpenStudy (anonymous):

The answer z B= -1.125

OpenStudy (anonymous):

Reciprocal is when the numerator and denominator are flipped. So,\[\frac{ -9b }{ 6b + 2 } \] Turns into... \[\frac{ 6b + 2 }{-9b }\] You need to simplify it though, right?

OpenStudy (anonymous):

ya ok thanks

terenzreignz (terenzreignz):

@jazy simplify or not, you end up with a fraction bar anyway... maybe it's best to leave it be?

OpenStudy (anonymous):

Agreed. (:

terenzreignz (terenzreignz):

Well, as I'm here anyway, the reciprocal r of a fraction \[\frac{p}{q}\] is whatever you have to multiply to it to get 1, so... \[ (r ) \ \frac{p}{q} = 1 \] or \[\huge r = \frac{1}{\frac{p}{q}}\]\[ r = \frac{q}{p}\] Or, as was said, just flip over the numerator and denominator :)